OpenStudy (anonymous):

equation (a²+b²)x²-2(ac+bd)x+(c²+d²)=0 are equal, prove that a/b=c/d.

myininaya (myininaya):

Using the Quadratic Formula, we obtain: \[x=\frac{2(ac+bd) \pm \sqrt{(-2(ac+bd))^2-4(a^2+b^2)(c^2+d^2)}}{2(a^2+b^2)}\] but I'm not sure how we can show a/b=c/d from this

myininaya (myininaya):

why did you say are equal? that's what the = sign signs say

myininaya (myininaya):

so if we can show ad-bc=0 we can show a/b=c/d so if we can ad-bc is a zero then we win

OpenStudy (anonymous):

bca mathematics this ,how to solve i don't yar plz help me

myininaya (myininaya):

i was just asking because i thought you may be missing something since you have=0 are equal simplify that that i have above

myininaya (myininaya):

\[x=\frac{2(ac+bd) \pm \sqrt{4(ac+bd)^2-4(a^2c^2+a^2d^2+c^2b^2+b^2d^2)}}{2(a^2+b^2)}\] \[x=\frac{2(ac+bd) \pm \sqrt{4} \sqrt{(ac+bd)^2-(a^2c^2+a^2d^2+c^2b^2+b^2d^2)}}{2(a^2+b^2)}\] \[x=\frac{2(ac+bd) \pm 2 \sqrt{(ac+bd)(ac+bd)-a^2c^2-a^2d^2-c^2b^2-b^2d^2}}{2(a^2+b^2)}\] \[x=2 \cdot \frac{(ac+bd) \pm \sqrt{a^2c^2+acbd+acbd+b^2d^2-a^2c^2-a^2d^2-c^2b^2-b^2d^2}}{2(a^2+b^2)}\] \[x=\frac{ac+bd \pm \sqrt{2acbd-a^2d^2-c^2b^2}}{a^2+b^2}\] \[x=\frac{ac+bd \pm \sqrt{-(a^2d^2-2acbd+c^2b^2)}}{a^2+b^2}\] \[x=\frac{ac+bd \pm \sqrt{-(ad-cb)^2}}{a^2+b^2}\] \[x=\frac{ac+bd \pm \sqrt{(ad-cb)^2} \sqrt{-1}}{a^2+b^2}\] \[x=\frac{ac+bd \pm (ad-cb)i}{a^2+b^2}\] ....

myininaya (myininaya):

yeah i don't think we can show a/b=c/d unless your teacher said such that the equation only has real root(s)

myininaya (myininaya):

if she said that part then ad-cb would have to =0 in order for us to have ad-cb=0 ad=cb a/b=c/d

OpenStudy (jamesj):

right. If the roots of the equation are equal, then ....

myininaya (myininaya):

i was trying to figure out why he said =0 are equal

myininaya (myininaya):

like there is suppose to be something there?

OpenStudy (jamesj):

definitely

OpenStudy (jamesj):

So if we wrote this equation as Ax^2 + Bx + C = 0 the roots are equal if B^2 - 4AC = 0 Now A = a^2 + b^2 etc.

OpenStudy (jamesj):

Then substitute and grind away a little

myininaya (myininaya):

right is that what you think he means by are equal?

OpenStudy (jamesj):

yes, definitely

myininaya (myininaya):

ok i was having problem with why would someone say something is equal twice? a=b are equal lol

myininaya (myininaya):

= means equal so why you say a=b are = its just weird

myininaya (myininaya):

thanks James

OpenStudy (jamesj):

So, let's just check for a minute: A = a²+b² B = -2(ac+bd) C = c²+d² Then B^2 - 4AC = 4[ (ac+bd)^2 - (a^2 + b^2)(c^2 + d^2) ] = 0 if and only if 2abcd = (ad)^2 + (bc)^2 i.e., (ad - bc)^2 = 0 i.e., a/b = c/d